OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8)? show work pls

OpenStudy (anonymous):

A polynomial with roots a and b has an equation:y=k(x-a)(x-b) here we have y=k[x-(-10)][x-(-6)]=k(x+10)(x+6) now we use our given point: (-8,-8) if we let x=-8, we get y=k(-8+10)(-8+6)=k(2)(-2)=-4k -8=-4k k=2 that means the equation is y=2(x+10)(x+6) is this correct?^^

OpenStudy (anonymous):

That's correct :) I made an error in my work, you're right